OpenStudy (anonymous):

solve for x: 1/2x+6=3+2x

OpenStudy (anonymous):

1/2x+6=3+2x [subtract 3 from both] 1/2x+3=2x [subtract 1/2x from both] 3=[3/2]x [divide both by 3/2] x = 2